Data:
length of pendulum =l=0.5m
acceleration =g=9.8m/s²
To find:
frequency =?
1
Expert's answer
2020-06-11T10:35:24-0400

Given data,

Length of pendulum is "l=0.5m"

Acceleration due to gravity "g=9.8m\/s^2"


Since,

"T=2\pi\sqrt{\frac{l}{g}}\\
\implies \nu=\frac{1}{T}=\frac{1}{2\pi}\sqrt{\frac{g}{l}}"

Thus,

"\nu=\frac{1}{2\pi}\sqrt{\frac{9.8}{0.5}}=0.70Hz"